OpenStudy (anonymous):

integration (x/sqrt(27+6x-x^2) dx = ?

OpenStudy (cherylim23):

Hint, you will need to use trig identity d/dx sin(x) = 1/sqrt(1-x^2) and linear substitution here.

OpenStudy (anonymous):

can u tell me, what should i do as the first step?

OpenStudy (cherylim23):

int_ [x/sqrt(36-(x-3)^2)] dx Let u =36 - (x-3)^2, du/dx = -2x +6 Keep going from there

OpenStudy (ash2326):

@tanjung do you get this?

OpenStudy (anonymous):

wait still process ^^

OpenStudy (anonymous):

i get this answer is - √(36-(x-3)^2 ) + 3 arc(sin(x-3)/6) + c, right?? sorry, i can't show my job.. because i dont know how to inputs (type) intgrl symbol here :(

OpenStudy (cherylim23):

Looks right to me

OpenStudy (anonymous):

i want to know, how to put the symbols of math here.. please tell me, sorry i dont understand yet